# Simulating conditioned diffusions on manifolds

### v2

@@ Abstract.
To date, most methods for simulating conditioned diffusions are limited to the Euclidean setting. The conditioned process can be constructed using a change of measure known as Doob’s $h$-transform. The specific type of conditioning depends on a function $h$ which is typically unknown in closed form. To resolve this, we extend the notion of guided processes to a manifold $M$, where one replaces $h$ by a function based on the heat kernel on $M$. We consider the case of a Brownian motion with drift, constructed using the frame bundle of $M$, conditioned to hit a point $x_{T}$ at time $T$. We prove equivalence of the laws of the conditioned process and the guided process with a tractable Radon-Nikodym derivative. Subsequently, we show how one can obtain guided processes on any manifold $N$ that is diffeomorphic to $M$ without assuming knowledge of the heat kernel on $N$. We illustrate our results with numerical simulations and an example of parameter estimation where a diffusion process on the torus is observed discretely in time.
Keywords: bridge simulation, Doob’s $h$-transform, geometric statistics, guided processes, Riemannian manifolds
@@ 1.4. Outline
In Section 2, we give some background on SDEs on manifolds. All main results of the paper are stated in Sections 3 and 4. We detail the approach taken in this paper in Section 3 in the specific case that $\Phi$ is the identity. In Section 4 we extend this result to a case where $\Phi$ is a diffeomorphism between manifolds $M$ and $N$ and the heat kernel is known on $N$. In Section 5, we illustrate the usefulness of our results in numerical simulations of diffusion bridges and a parameter estimation problem for a discretely observed diffusion process on the torus. In Section 6, we show how our approach can be extended to other types of conditionings at a fixed future time. Finally, all proofs are gathered in Section 7.
@@ 2. Background on SDEs on manifolds
We consider SDEs on manifolds as developments of ${\mathbb{R}}^{d}$-valued semimartingales. We sketch the main concepts here. For a detailed description, we refer to Hsu, (2002); Emery, (1989). With $M$ a $d$-dimensional Riemannian manifold and $x\in M$, the tangent space $T_{x}M$ to $M$ at $x$ is a $d$-dimensional vector space. A frame $u$ is a basis for $T_{x}M$, equivalently an isomorphism $u\colon{\mathbb{R}}^{d}\to T_{x}M$. The collection of all frames at $x$ is denoted by $\mathrm{F}_{x}M$ and the vector bundle with fibers $\mathrm{F}_{x}M$ is called the frame bundle $\mathrm{F}M$. The canonical projection, which maps a frame $u$ at $x$ to $x$, is denoted by $\pi\colon{\mathrm{F}M}\to M$.
${\mathrm{F}M}$ is a manifold itself of dimension $d+d^{2}$. It has a submanifold ${\mathrm{O}M}$ consisting of orthonormal frames. A derivative of a curve in ${\mathrm{F}M}$ can be decomposed in a horizontal part, which corresponds to infinitesimal parallel translation of the basis along the projection onto $M$, and a vertical part, which is derivative of a curve for which the projection stays constant but the frame changes over time. Therefore, for $u\in{\mathrm{F}M}$, $T_{u}{\mathrm{F}M}$ can be written as a direct sum $T_{u}{\mathrm{F}M}=\mathcal{H}_{u}{\mathrm{F}M}\oplus\mathcal{V}_{u}{\mathrm{F}M}$ of the horizontal and vertical subspaces, respectively. The horizontal subspace is a vector space of dimension $d$ with a basis $H_{1}(u),\dots,H_{d}(u)$ consisting of the globally defined horizontal fields $H_{1},\ldots,H_{d}$ that in local coordinates at $u=(x^{i},\nu^{i}_{j})$ are given by
$H_{i}(u)=\nu^{j}_{i}\frac{\partial}{\partial x^{j}}-\nu^{j}_{i}\nu^{l}_{m}\Gamma^{k}_{jl}\frac{\partial}{\partial\nu^{k}_{m}}$
An ${\mathbb{R}}^{d}$-valued semimartingale $Z$ can be uniquely mapped to a horizontal process $U$ via the Stratonovich SDE
@@
The process $U$ is called the development of $Z$ and the projection $\pi(U)$ is the development of $Z$ on $M$. The development is reversible in the sense that, for a horizontal semimartingale $U$, a unique ${\mathbb{R}}^{d}$-valued semimartingale $Z$ exists so that $U$ is the development of $Z$. This $Z$ is denoted the anti-development of $U$. It can be shown that if $W$ is an ${\mathbb{R}}^{d}$-valued Brownian motion and $U_{0}=u_{0}\in{\mathrm{O}M}$, then $\pi(U)$ is Brownian motion in $M$.
If $X$ is a semimartingale on $M$, a horizontal semimartingale $U$ such that $\pi(U)=X$ is denoted a horizontal lift of $X$. For a function $f\colon M\to{\mathbb{R}}$, we denote by $\tilde{f}\colon{\mathrm{F}M}\to{\mathbb{R}}$ the function $f\circ\pi$. For $u\in{\mathrm{F}M}$, a vector $V\in T_{\pi(u)}M$ can be uniquely lifted to $\mathcal{H}_{u}{\mathrm{F}M}$ using the horizontal vector fields. This vector is then denoted by $\tilde{V}$. In particular, for a function $f\colon M\to{\mathbb{R}}$, we denote the lift of $\nabla f$ by $\nabla^{H}\tilde{f}=\left\{H_{i}\tilde{f},\dots,H_{d}\tilde{f}\right\}$ and refer to it as the horizontal gradient.
@@ 3. Guided processes on manifolds
where we again used Einsteins summation convention. Alternatively, one could also write $U$ under ${\mathbb{P}}^{h}$ as the stochastic development of an ${\mathbb{R}}^{d}$-valued process through
@@ 3.1. Guiding induced by the heat kernel
It follows from Sturm, (1993) that this assumption is satisfied if $M$ has Ricci curvature bounded from below, which is the case for compact manifolds. ∎
@@
Assume $M$ is a compact manifold and $V$ is a smooth vector field on $M$. Then ${\mathbb{P}}_{T}^{h}\sim{\mathbb{P}}_{T}^{g}$ with
@@ 4. Guiding with comparison manifolds
In view of (4.1), this is what one should expect, as $r(t,x)$ is a vector in $T_{x}{\mathbb{R}}^{d}$ and the new guiding term is $\Phi_{*}\left(\nabla_{x}\log h\right)(t,\Phi^{-1}(y))$ is exactly the pushforward of this tangent vector under $\Phi$.
@@ 5.1. Bridge simulation
Suppose we wish to sample bridges on the manifold $N$ connecting $y_{0}$ at time 00 and $y_{T}$ at time $T$. Suppose $\Phi\colon M\to N$ is a diffeomorphism and on $M$ the heat kernel is known in closed form. Assume $g$ as in (3.7) with $x_{T}=\Phi^{-1}(y_{T})$ and let $x_{0}=\Phi^{-1}(y_{0})$. Let $u_{0}$ be such that $x_{0}=\pi(u_{0})$ Assuming a strong solution to Equation 3.4, there exists a map $\mathcal{GP}$ such that $U=\mathcal{GP}(u_{0},Z)$. In our numerical implementation, this means that the SDE for $U$ is solved by Euler-Heun discretisation on a dense (fixed) grid, and $Z$ is the vector with increments of the driving Wiener process of the SDE on this grid. With slight abuse of notation we write $X=\pi(U)$ for the projection of $U$ on the manifold $M$.
Algorithm 1 gives a Metropolis-Hastings algorithm to sample from ${\mathbb{P}}^{h}$ on $N$. With $\lambda=0$ it corresponds to an Metropolis-Hastings algorithm with independent proposal distribution. Note that the guided process does take the drift $V$ into account, but ignores it in the guiding term. If the drift is strong, or we condition the process on a very unlikely point, the acceptance rate of the independence sampler may be very low. In that case, taking a value of $\lambda$ closer to $1$, for example to target a Metropolis-Hastings acceptance rate of about $25\%-50\%$, will facilitate the chain to reach stationarity. In the numerical simulation results that we present, $\lambda$ was chosen by monitoring the acceptance rates on a trial run and subsequently adjusted to achieve the desired acceptance rate. Alternatively, adaptive methods can be used for tuning $\lambda$ (cf. Roberts & Rosenthal, (2009)). In all simulations below we sampled the guided process on a grid with mesh width $0.001$, mapped through a time change $t\mapsto t(1.7-0.7t/T)$ to allow for more grid points near $T$.
@@
1 Sample a Brownian $Z$, set $U=\mathcal{GP}(x_{0},Z)$ and $X=\pi(U)$ ;
@@ 2 Sample a Brownian motion $W$ independent of $Z$ and let $Z^{\circ}=\lambda Z+\sqrt{1-\lambda^{2}}W$ ;
3 Simulate $U^{\circ}=\mathcal{GP}(x_{0},Z^{\circ})$, set $X^{\circ}=\pi(U^{\circ})$ and compute $\alpha=1\wedge\Psi_{T,g}(X^{\circ})/\Psi_{T,g}(X)$ with $\Psi_{T,g}$ defined in (3.9) ;
@@ 5 Repeat steps 2–4 $K$ times to obtain $X^{(1)},\ldots,X^{(K)}$;
$\kappa(t,x,y)=\frac{1}{2\pi t}\sum_{k\in{\mathbb{Z}}}\sum_{\ell\in{\mathbb{Z}}}\exp\left(-\frac{x^{1}-y^{1}-2k\pi}{2t}-\frac{x^{2}-y^{2}-2\ell\pi}{2t}\right).$ (5.1)
@@
    A somewhat weak vector field with $\theta=(-2,2)$, where we condition the process on a point obtained from forward simulation of the unconditional process (Figure 4, guided processes obtained with $\lambda=0.95$). One can visually assess that samples of the guided process the original forward path.
@@
5.2. Parameter estimation
Here, we perform inference on the parameter $\theta$ in (5.2) based on observations $\mathcal{D}=\{X_{t_{i}}=x_{i},\,i=1,\dots,n\}$. To perform parameter inference conditional on $\mathcal{D}$, we simultaneously sample $\theta,X\mid\mathcal{D}$. This can be done using Gibbs-type sampling method in which we repeatedly perform the following steps
  1. (1)
    Update $Z\mid(\mathcal{D},\theta)$ using steps 2–4 of Algorithm 1 and set $U=\mathcal{GP}_{\theta}(x_{0},Z)$.
  2. (2)
    Update $\theta\mid(\mathcal{D},X)$, where $X=\pi(U)$.
  3. (3)
    Compute $Z$ as the anti-development of $X$ under the drift with parameter $\theta$.
@@ 5.2. Parameter estimation
We applied this Gibbs sampler to a dataset generated under forward simulating $X$ with vector field (5.2) with $\theta=(4,-4)$ and saving its values at $40$ equidistant times. The posterior samples are displayed in Figure 7. For the prior, we used $\Gamma_{0}=0.05I$ and we used tuning parameter $\lambda=0.9$ in the Crank-Nicolson steps.
@@
Figure 7. Left: Histograms 1900 samples from the posterior distribution of $\theta$ after 100 iteration of burn-in time using the Gibbs sampler Right: Trace of the sampler. The top row are the samples for $\theta_{1}$, the bottom row the samples for $\theta_{2}$. Data were generated with $(\theta_{1},\theta_{2})=(4,-4)$ as indicated by the red lines.

### v3

@@ Abstract.
To date, most methods for simulating conditioned diffusions are limited to the Euclidean setting. The conditioned process can be constructed using a change of measure known as Doob’s $h$-transform. The specific type of conditioning depends on a function $h$ which is typically unknown in closed form. To resolve this, we extend the notion of guided processes to a manifold $M$, where one replaces $h$ by a function based on the heat kernel on $M$. We consider the case of a Brownian motion with drift, constructed using the frame bundle of $M$, conditioned to hit a point $x_{T}$ at time $T$. We prove equivalence of the laws of the conditioned process and the guided process with a tractable Radon-Nikodym derivative. Subsequently, we show how one can obtain guided processes on any manifold $N$ that is diffeomorphic to $M$ without assuming knowledge of the heat kernel on $N$. We illustrate our results with numerical simulations of guided processes and Bayesian parameter estimation based on discrete-time observations. For this, we consider both the torus and the Poincaré disk.
Keywords: bridge simulation, Doob’s $h$-transform, geometric statistics, guided processes, Poincaré disk, Riemannian manifolds
@@ 1.4. Outline
In Section 2, we give some background on SDEs on manifolds. All main results of the paper are stated in Sections 3 and 4. We detail the approach taken in this paper in Section 3 in the specific case that $\Phi$ is the identity. In Section 4 we extend this result to a case where $\Phi$ is a diffeomorphism between manifolds $M$ and $N$ and the heat kernel is known on $N$. In Section 5, we illustrate the usefulness of our results in numerical simulations of diffusion bridges and a parameter estimation problem for a discretely observed diffusion process on the torus and the Poincaré disk. In Section 6, we show how our approach can be extended to other types of conditionings at a fixed future time. Finally, all proofs are gathered in Section 7.
@@ 2. Background on SDEs on manifolds
We consider SDEs on manifolds as developments of ${\mathbb{R}}^{d}$-valued semimartingales. We sketch the main concepts here. For a detailed description, we refer to Hsu, (2002); Emery, (1989). With $M$ a $d$-dimensional Riemannian manifold and $x\in M$, the tangent space $T_{x}M$ to $M$ at $x$ is a $d$-dimensional vector space. A frame $u$ is a basis for $T_{x}M$, which can be seen as an isomorphism $u\colon{\mathbb{R}}^{d}\to T_{x}M$ where an ${\mathbb{R}}^{d}$-valued vector is mapped to $T_{x}M$ as a linear combination of the basis vectors in $u$. The collection of all frames at $x$ is denoted by $\mathrm{F}_{x}M$ and the vector bundle with fibers $\mathrm{F}_{x}M$ is called the frame bundle $\mathrm{F}M$. The canonical projection, which maps a frame $u$ at $x$ to $x$, is denoted by $\pi\colon{\mathrm{F}M}\to M$. In short, each element of ${\mathrm{F}M}$ can be interpreted as a tuple $u=(x,\nu)$ with $x\in M$ and $\nu=(\nu_{1},\dots,\nu_{d})$ where $\nu_{1},\dots,\nu_{d}\in T_{x}M$ form a basis for the vector space $T_{x}M$. The canonical projection $\pi$ is simply the map $(x,\nu)\mapsto x$. The development ${\mathbb{R}}^{d}$-semimartingales onto $M$ refers to a a map that first maps the semimartingale to $\mathrm{F}M$ and then projects it onto $M$. The reason why this is convenient is that the increment $dZ_{t}$ of an ${\mathbb{R}}^{d}$-semimartingale is in ${\mathbb{R}}^{d}$ and can thus, through a frame, be mapped to the tangent space of $M$, which then gives a direction for the $M$-valued process to move to.
It can be shown that ${\mathrm{F}M}$ is a manifold itself of dimension $d+d^{2}$. It has a submanifold ${\mathrm{O}M}$ consisting of all the orthonormal frames, i.e. frames $(x,\nu)$ of which the vectors in $\nu$ are orthonormal with respect to the Riemannian inner product in $T_{x}M$. Given $u\in\mathrm{F}M$, we have to study the tangent space $T_{u}{\mathrm{F}M}$ in order to make sense of stochastic increments. Since the tangent space of a manifolds consists of derivatives of curves on the manifold, we consider ${\mathrm{F}M}$-valued curves first. A curve in ${\mathrm{F}M}$ can be decomposed in a horizontal part, which corresponds parallel transport of the basis along the projection onto $M$, and a vertical part, which is the part where the projection stays constant but only the basis changes over time. This decomposition extends to the derivative of a curve and therefore, for $u\in{\mathrm{F}M}$, $T_{u}{\mathrm{F}M}$ can be written as a direct sum $T_{u}{\mathrm{F}M}=\mathcal{H}_{u}{\mathrm{F}M}\oplus\mathcal{V}_{u}{\mathrm{F}M}$ of the horizontal and vertical subspaces, respectively. The projection ${\mathrm{F}M}\to M$ can be extended to a projection $T_{u}{\mathrm{F}M}\to T_{\pi u}M$, but what is interesting is that, when restricting ourselves to $\mathcal{H}_{u}{\mathrm{F}M}$, this projection is bijective, and can thus be inverted to a “lift” $T_{x}M\to T_{u}{\mathrm{F}M}$, where $u$ is a frame with $\pi u=x$. It is therefore very useful to look at the horizontal subspace, as the vertical part of the process is irrelevant with us mainly being interested in the projection onto $M$. The horizontal subspace is a vector space of dimension $d$ with a basis $H_{1}(u),\dots,H_{d}(u)$ consisting of the globally defined “horizontal fields” $H_{1},\ldots,H_{d}$ that in local coordinates at $u=(x^{i},\nu^{i}_{j})$ are given by
$H_{i}(u)=\nu^{j}_{i}\frac{\partial}{\partial x^{j}}-\nu^{j}_{i}\nu^{l}_{m}\Gamma^{k}_{jl}\frac{\partial}{\partial\nu^{k}_{m}},$
where $\Gamma^{k}_{jl}$ denote Christoffel symbols on $M$. An ${\mathbb{R}}^{d}$-valued semimartingale $Z$ can be uniquely mapped to a horizontal process $U$ via the Stratonovich SDE
@@
The process $U$ is called the development of $Z$ and the projection $\pi(U)$ is the development of $Z$ on $M$. The development is invertible in the sense that, for a horizontal semimartingale $U$, a unique ${\mathbb{R}}^{d}$-valued semimartingale $Z$ exists so that $U$ is the development of $Z$. This $Z$ is denoted the anti-development of $U$. It can be shown that if $W$ is an ${\mathbb{R}}^{d}$-valued Brownian motion and $U_{0}=u_{0}\in{\mathrm{O}M}$, then $\pi(U)$ is Brownian motion in $M$.
If $X$ is a semimartingale on $M$, a horizontal semimartingale $U$ such that $\pi(U)=X$ is referred to as a horizontal lift of $X$. For a function $f\colon M\to{\mathbb{R}}$, we denote by $\tilde{f}\colon{\mathrm{F}M}\to{\mathbb{R}}$ the function $f\circ\pi$. For $u\in{\mathrm{F}M}$, a vector $V\in T_{\pi(u)}M$ can be uniquely lifted to $\mathcal{H}_{u}{\mathrm{F}M}$ using the horizontal vector fields. This vector is then denoted by $\tilde{V}$. In particular, for a function $f\colon M\to{\mathbb{R}}$, we denote the lift of $\nabla f$ by $\nabla^{H}\tilde{f}=\left\{H_{i}\tilde{f},\dots,H_{d}\tilde{f}\right\}$ and refer to it as the horizontal gradient.
@@ 3. Guided processes on manifolds
where we again used Einstein’s summation convention. Alternatively, one could also write $U$ under ${\mathbb{P}}^{h}$ as the stochastic development of an ${\mathbb{R}}^{d}$-valued process through
@@ 3.1. Guiding induced by the heat kernel
This result follows from Equation (1.2) of Sturm, (1993) if $M$ has Ricci curvature bounded from below, which is the case for compact manifolds. ∎
@@
Assume $M$ is a compact manifold and $V$ is a smooth vector field on $M$. Let $h\in\mathcal{H}$ be such that ${\mathfrak{A}}h=0$ and let $g$ be as in (3.7). Then ${\mathbb{P}}_{T}^{h}\sim{\mathbb{P}}_{T}^{g}$ with
@@ 4. Guiding with comparison manifolds
In view of (4.1), this is what one should expect, as $r(t,x)$ is a vector in $T_{x}{\mathbb{R}}^{d}$ and the new guiding term $\Phi_{*}\left(\nabla_{x}\log h\right)(t,\Phi^{-1}(y))$ is exactly the pushforward of this tangent vector under $\Phi$.
@@ 5.1. Bridge simulation
Suppose we wish to sample bridges on the manifold $N$ connecting $y_{0}$ at time 00 and $y_{T}$ at time $T$. Suppose $\Phi\colon M\to N$ is a diffeomorphism and on $M$ the heat kernel is known in closed form. Assume $g$ as in (3.7) with $x_{T}=\Phi^{-1}(y_{T})$ and let $x_{0}=\Phi^{-1}(y_{0})$. Let $u_{0}$ be such that $x_{0}=\pi(u_{0})$. Assuming a strong solution to Equation 3.4, there exists a map $\mathcal{GP}$ such that $U=\mathcal{GP}(u_{0},Z)$. In our numerical implementation, this means that the SDE for $U$ is solved by Euler-Heun discretisation on a dense (fixed) grid, and $Z$ is the vector with increments of the driving Wiener process of the SDE on this grid. With slight abuse of notation we write $X=\pi(U)$ for the projection of $U$ on the manifold $M$.
Algorithm 1 gives a Metropolis-Hastings algorithm to sample from ${\mathbb{P}}^{h}$ on $N$ using a preconditioned Crank-Nicolson step (Cotter2013MCMC). With $\lambda=0$ it corresponds to an Metropolis-Hastings algorithm with independent proposal distribution. Note that the guided process does take the drift $V$ into account, but ignores it in the guiding term. If the drift is strong, or we condition the process on a very unlikely point, the acceptance rate of the independence sampler may be very low. In that case, taking a value of $\lambda$ closer to $1$, for example to target a Metropolis-Hastings acceptance percentage of about $25\%-50\%$, will facilitate the chain to reach stationarity. In the numerical simulation results that we present, $\lambda$ was chosen by monitoring the acceptance rates on a trial run and subsequently adjusted to achieve the desired acceptance rate. Alternatively, adaptive methods can be used for tuning $\lambda$ (cf. Roberts & Rosenthal, (2009)). In the simulation results on the torus, we sampled the guided process on a grid with mesh width $0.001$, mapped through a time change $t\mapsto t(1.7-0.7t/T)$ to allow for more grid points near $T$.
@@
1 Sample a Brownian motion $Z$, set $U=\mathcal{GP}(u_{0},Z)$ and $X=\pi(U)$ ;
@@ 2 Sample a Brownian motion $W$ independent of $Z$ and let $Z^{\circ}=\lambda Z+\sqrt{1-\lambda^{2}}W$ ;
3 Set $U^{\circ}=\mathcal{GP}(u_{0},Z^{\circ})$, set $X^{\circ}=\pi(U^{\circ})$ and compute $\alpha=1\wedge\Psi_{T,g}(X^{\circ})/\Psi_{T,g}(X)$ with $\Psi_{T,g}$ defined in (3.9) ;
@@ 5 Repeat steps 2–4 $K$ times to obtain $X^{(1)},\ldots,X^{(K)}$;
5.2. Parameter estimation
Suppose we have observations $\mathcal{D}=\{X_{t_{i}}=x_{i},\,i=1,\dots,n\}$ and wish to infer the parameter $\theta$ in (5.2). To perform parameter inference conditional on $\mathcal{D}$, we use data augmentation, where we construct a Markov chain that samples from the joint distribution of $(\theta,X)$, conditional on $\mathcal{D}$. Here, $X=\{X_{t},\,t\in[0,t_{n}]\}$. Specifically, we repeatedly perform the following steps
  1. (1)
    Update $Z\mid(\mathcal{D},\theta)$ using steps 2–4 of Algorithm 1 and set $U=\mathcal{GP}_{\theta}(u_{0},Z)$.
  2. (2)
    Update $\theta\mid(\mathcal{D},X)$, where $X=\pi(U)$.
  3. (3)
    Compute $Z$ as the anti-development of $X$ under the drift with parameter $\theta$. That is,
    $Z_{t}=\int_{0}^{t}U_{s}^{-1}\circ\left(\,\mathrm{d}X_{s}-V_{\theta}(X_{s})\,\mathrm{d}s-\nabla\log g(s,X_{s})\,\mathrm{d}s\right)$
5.3. Example: guided processes on the torus
@@
$\kappa(t,x,y)=\frac{1}{2\pi t}\sum_{k\in{\mathbb{Z}}}\sum_{\ell\in{\mathbb{Z}}}\exp\left(-\frac{(x_{1}-y_{1}-2k\pi)^{2}}{2t}-\frac{(x_{2}-y_{2}-2\ell\pi)^{2}}{2t}\right).$ (5.1)
@@
    A somewhat weak vector field with $\theta=(-2,2)$, where we condition the process on a point obtained from forward simulation of the unconditional process (Figure 4, guided processes obtained with $\lambda=0.95$). One can visually assess that samples of the guided process resemble the original forward path.
@@
We applied the algorithm for parameter estimation from Section 5.2 to a dataset generated under forward simulating $X$ with vector field (5.2) with $\theta=(4,-4)$ and saving its values at $40$ equidistant times. The posterior samples are displayed in Figure 7. For the prior, we used $\Gamma_{0}=0.05I$ and we used tuning parameter $\lambda=0.9$ in the Crank-Nicolson step.
@@
Figure 7. Left: Histograms 1900 samples from the posterior distribution of $\theta$ after 100 iterations of burn-in time using the Gibbs sampler Right: Trace of the sampler. The top row are the samples for $\theta_{1}$, the bottom row the samples for $\theta_{2}$. Data were generated with $(\theta_{1},\theta_{2})=(4,-4)$ as indicated by the red lines.
5.4. Example: guided processes on the Poincaré disk
Hansen et al., (2021) provide several examples of manifolds on which the heat kernel is known. On all manifolds diffeomorphic to these, the approach presented in this article can be applied. Since diffusion models on hyperbolic spaces are often seen in machine learning applications, see e.g. fu2024hyperbolic, wen2024hyperbolicgraphdiffusionmodel as well as in the theory of general relativity, we’ll explore this setting in more detail here.
The hyperbolic space is a non-Euclidean geometric space with constant negative sectional curvature. Unlike Euclidean spaces, where parallel lines remain equidistant, in hyperbolic space, they diverge. The $2$-dimensional hyperpolic space can be embedded into t ${\mathbb{R}}^{3}$, equipped with the Minkowski metric by considering the subspace of all points with Minkowski norm $-1$. Here, we consider the Poincaré disk to model the $2$-dimensional hyperbolic space in local coordinates, which can be interpreted as a stereographic projection of the three-dimensional representation onto the unit disk. Here the entire infinite space is mapped inside the unit disk, preserving angles but distorting distances. Let $D=\{z\in\mathbb{C}\,:\,|z|<1\}$. The hyperbolic distance between $z_{1},z_{2}\in D$ is given by
${\mathrm{dist}}(z_{1},z_{2})=\operatorname{arcosh}\left(1+\frac{2|z_{1}-z_{2}|^{2}}{(1-|z_{1}|^{2})(1-|z_{2}|^{2})}\right)$
An expression for the heat kernel for general $d$ on $\mathcal{H}^{d}$ is given in grigoryan_hyperbolic. For $d=2$, we have
$p(s,x;t,y)=\frac{\sqrt{2}e^{-(t-s)/2}}{(2\pi(t-s))^{3/2}}\int_{\rho_{x}}^{\infty}\frac{u\exp\left(-\frac{u^{2}}{2(t-s)}\right)}{\sqrt{\cosh u-\cosh\rho_{x}}}\,\mathrm{d}u,$ (5.3)
where $\rho_{x}={\mathrm{dist}}(x,y)$ denotes the geodesic distance between $x$ and $y$.
Despite the manifold not being compact, we apply our methodology and construct the guided process. Details on evaluating $\nabla_{x}\log p(s,x;t,y)$ are given in Appendix B. We consider three vector fields
  • $V_{1}(x)=-20x$ which strongly direct paths towards the center of the disc;
  • $V_{2}(x)=x$ which sends paths towards the boundary of the disc;
  • $V_{3}(x)=[5(1-x_{1}^{2}-x_{2}^{2})^{2};0]$ which sends paths in the east direction. A visualisation of this vector field is in the left panel of Figure 9.
For each of these vector fields, we simulate $4$ independent realisations of the guided process on $[0,1]$ where each path starts at $x_{0}=[0;0.6]$ and ends at $x_{1}=[0;-06]$. The paths were simulated on a grid with mesh-width $0.001$ that was then mapped to a non-equidistant grid through the map $s\mapsto s(2-s)$ as advocated in Section 5 of meulen_schauer_bayest. The simulated paths are shown in Figure 8. One can see in the top panel that the process is pushed rather quickly towards the center, stays most of the time near the center, and then moves to $x_{1}$. The bottom panel indeed show the paths tend to go eastwards.
Figure 8. Each panel shows an independent realisation of the guided process connecting $x_{0}$ (red dot) and $x_{1}$ (blue dot). Top row: vector field $V_{1}$, middle row: vector field $V_{2}$; bottom row: vector field $V_{3}$.
Figure 9. Left: visualisation of vector field $V_{3}$. Right: each panel shows an independent realisation of the guided process connecting $x_{0}=[0;0.2]$ (red dot) and $x_{1}=[-0.8;0]$ (blue dot).
In the right panel of Figure 9 we show two simulated guided paths where we changed the starting point to $[0;0.2]$ and condition the process to hit $x_{1}=[-0.8;0]$ at time $1$. One can see that paths initially tend to move eastwards due to the vector field $V_{3}$ but are subsequently forced to hit $x_{1}$ at time $1$.
For vector field $V_{3}$ we run the pCN-scheme of Algorithm 1 with $\lambda=0.4$ for $500$ iterations and show every $50$th iterate in Figure 10. The acceptance percentage was approximately $80\%$, showing that the guided process qualitatively approximates the true conditioned process quite well.
Figure 10. Every $50$th iteration out of $500$ iterations after running the pCN-scheme of Algorithm 1 for vector field $V_{3}$ on $\mathcal{H}^{2}$.
Lastly, we apply the Gibbs sampler described Section 5.2 to estimate the parameter $\theta$ appearing in the vector field $V_{\theta}(x)=[\theta(1-x_{1}^{2}-x_{2}^{2})^{2};0]$ with a $N(0,20^{2})$-prior. Figure 11 demonstrates the estimates of 1000 iterations with tuning parameter $\lambda=0.9$. The observations were observations of a forward simulated process with $\theta=5$ on an equally spaced grid of 100 points in the time interval $[0,1]$.
Figure 11. Top: Histogram of 900 from the posterior distribution of the parameter $\theta$ appearing in the vector field $V_{\theta}(x)=[\theta(1-x_{1}^{2}-x_{2}^{2})^{2};0]$ after omitting 100 burn-in samples. Below: Trace of the sampler. The samples are based on 100 observations of a forward simulated process with $\theta=5$
These simulations suggest that the results of this article may also hold on the Poincaré disk, which is an example of a noncompact manifold.